### v1

# Capsule: Efficient Player Isolation for Datacenters

Abstract.
Cloud gaming is increasingly popular. A challenge for cloud provider is to keep datacenter utilization high: a non-trivial task due to application variety. These applications come in different shapes and sizes. So do cloud datacenter resources, e.g., CPUs, GPUs, NPUs.
Part of the challenge stems from game engines being predominantly designed to run only one player. One player in a lightweight game might utilize only a fraction of the cloud server GPU. The remaining GPU capacity will be left underutilized, an undesired outcome for the cloud provider. We introduce Capsule, a mechanism that allows multiple players to seamlessly share one GPU.
We implemented Capsule in O3DE, a popular open source game engine. Our evaluations show that Capsule can increase datacenter resource utilization by accommodating up to 2.25x more players, without degrading player gaming experience. Capsule is also application agnostic. We ran four applications on Capsule-based O3DE with no application changes. Our experiences show that Capsule design can be adopted by other game engines to increase datacenter utilization across cloud providers.
copyright: none
1. Introduction
Cloud gaming is attractive for players as well as cloud providers. For players, it alleviates the deployment cost. They no longer have to own the latest hardware (e.g., GPU) to play the game in high quality. Cloud already hosts the latest hardware, sometimes before they become publicly available (Petty et al., 2023). For providers, it is about generating revenue while delivering the highest gaming quality. Higher the cloud datacenter utilization, higher the revenue.
However, it is challenging to achieve high datacenter utilization with gaming applications. Part of the challenge arises from games having diverse shapes and sizes. Shapes correspond to different resources games consume, such as CPU, RAM, GPU, and VRAM. Sizes correspond to their different amount, e.g., a graphics-intensive game consumes the entire GPU while a graphics-light game consumes only fraction of the GPU. Another challenge arises from datacenters having diverse hardware. A datacenter has servers with several generations of CPUs and GPUs. For example, a server with the oldest GPU could accommodate only one player while a server with the latest GPU accommodates dozens. Thus, cloud providers need a mechanism to share GPU, as well as other resources across multiple players.
Figure 2. Existing isolation mechanisms in the cloud. Lambdas are lightweight but require knowledge of application semantics. Virtual Machines are agnostic to semantics but are heavyweight. Container approach strikes the right balance for game engines.
Games are not the only kind of application that require cloud resource multiplexing. Other cloud workloads also suffer this problem and resource virtualization has been the primary solution for over 20 years (Barham et al., 2003). Broadly, there are three virtualization categories: Virtual Machines (VM), containers, and lambdas, as shown in Figure 2. The container approach is well-fit for gaming applications because (1) it does not require knowing the application semantics, which makes the solution applicable to wide range of games, and (2) it does not impose high overhead, i.e., multiplexing mechanism itself consumes insignificant CPU, RAM, and GPU resources. Lambdas suffer (1) and VMs suffer (2). How does container-like multiplexing solution look like for gaming applications?
We propose player-level multiplexing. A player in the graphics-heavy application will continue consuming the entire GPU. However, when a GPU has sufficient capacity to accommodate two or more players in a multiplayer game, GPU resources will be multiplexed across these players. We designed, implemented, and evaluated Capsule: an in-game-engine player isolation mechanism for multiplayer games. Capsule also allows cross-player sharing without cross-player performance inference. For example, when two players enter a room and have a shared game asset in their view, we can reuse the asset geometry across these two players, without players noticing. Sharing allows cloud providers to have a sublinear resource consumption growth for linear player increase–the phenomena we call super-linear player count, or superlinear for short.
We implemented Capsule in O3DE, a popular open source game engine (O3DE, 2025b). Our evaluations show that Capsule-based O3DE can increase datacenter resource utilization by accommodating up to 2.25x more players, without degrading player gaming experience. Capsule is also application agnostic. We ran four applications on Capsule with no application changes. Our experiences show that Capsule design is generalizable and can be adopted by other engines.
2. Design and Implementation
There are four requirements for a player-isolation mechanism in cloud, in order of their priority:
  • R1: Transparent: Players should be unaware of other players sharing cloud resources. A player experience, such as, streaming quality and frames-per-second (FPS), should not degrade due to other players.
  • R2: Compatible: Do not require significant changes to run existing applications, best if no application changes. The workflow for developing a new application should also remain near identical, if not exactly identical.
  • R3: Lightweight: Isolation should not consume significant system resources, such as CPU and RAM.
  • R4: Efficient: Maximize cross-player sharing, i.e., superlinearity. For example, resource footprint (e.g., RAM) of the second player should be less than that of the first player because of sharing.
Capsule satisfies all four requirements. However, for example, the process-level-isolation would violate R1, R3, and R4. The process-level-isolation is achieved when we simply run separate game engine process in a cloud server for each player. Here, R1 is violated because a player can consume more GPU resources during graphics-heavy scene, reducing the FPS for other players. R3 is violated because spawning and managing separate OS process is not lightweight. R4 is violated because players are unable to share computation results–a by-product of process-level-isolation by design.
However, Capsule introduces player fate-sharing, which might degrade player fault-tolerance. For example, when a server GPU fails, all players running on the game engine hosted on that GPU also fail. However, fate-sharing is common and inevitable in cloud: VMs, containers, lambdas all suffer it (Figure 2), e.g., when a server fails all VMs hosted on that server fail. The disadvantages of the fate-sharing can be ameliorated by other mechanisms, such as player migration with player state replication, as it has been done for VMs over 20 years ago (Clark et al., 2005).
Figure 3. Capsule-based cloud architecture. Capsule provides player isolation by relying on Capsule Storage for player state management.
Figure 3 shows Capsule architecture, along with other essential modules in cloud deployment. Capsule is a new module in O3DE. Capsule communicates with different system components, such as audio system, input system, rendering system (includes both audio and video rendering), and game logic (event system). Capsule leverages Entity–Component–System (ECS), an existing game engine architecture (Wikipedia, 2013). ECS makes it convenient to represent game world objects. An ECS-based game engine contains entities that have data components and systems to operate on those components. This architecture is widely adopted by modern game engines (Unity, 2025), including O3DE.
Capsule distinguishes different players’ input, output, and behaviours by using Capsule Storage. Capsule Storage has two layers: entities and events. Entities are game objects, such as a car, a light, an avatar. These are the key elements that decide what the final rendered video or audio is. Events, or gameplay events, include in-game events, such as running, jumping, exploding. They are predesigned by game developers and written inside game script. These are the key elements that decide the behavior of each player and the behaviour of global environment.
As shown in Figure 3, Capsule has two types of Storage: global and local. There is only one global storage in the entire engine. All players share the entities and events inside the global storage. There are one or more local Capsule Storages, one for each player. Per-player events are contained within the local storage. Capsule automatically isolates player-specific tasks at runtime by directing per-player entities and events to the respective player-owned local storage.
We ported four applications to Capsule-based O3DE. Figure 1 shows three of them for brevity. Our experience with these four applications show that Capsule-based O3DE is fully compatible (R2) with non-Capsule O3DE. No changes were required on these applications to run them on Capsule-based O3DE. Moreover, the workflow for further development of these applications on Capsule-based O3DE was identical to that of non-Capsule O3DE.
Capsule design principles can be easily adopted in other game engines thanks to ECS architecture. We also believe that four design requirements (R1R4) are the common goals in all cloud deployments. Thus, our design and implementation decisions are widely applicable to other game engines as well as other cloud environments.
3. Evaluation
We evaluated Capsule on a Windows workstation with AMD Ryzen 7 5800X 8-Core processor, 32 GB of RAM, and NVIDIA GeForce RTX 4090 GPU with 24 GB of VRAM. We fixed the application FPS to 30, a common minimum threshold. We compare Capsule resource footprints against the baseline that is implemented as process-level-isolation, i.e., a separate game engine process for each player. Note that this is an overly optimistic baseline for lightweightness comparison because it does not satisfy the transparency requirement (R1), as mentioned earlier. A realistic baseline that satisfies this requirement would inherently be less lightweight. Thus, the measured benefits of Capsule are conservative.
Figure 1 shows that Capsule can accommodate up to 8 players without dropping FPS below the threshold. Figure 1(a) and Figure 1(b) show game server view with single player (1p) and two players (2p), respectively. Figure 1(c) shows resource utilizations for these two environments. As shown in Figure 1(c), 2p GPU utilization (57%–67%) is up to 39% point higher than the baseline (28%–44%). This delta is smaller, but is more significant in other resources. If the delta is zero (1p and 2p lines overlap), it means 2p utilization of that resource is identical to that of 1p, which means the second player came for free (for that resource). In Figure 1(c), this is almost the case for CPU, RAM, and VRAM, i.e., the second player has a small extra cost. Note that Capsule is unable to accommodate more than two players because the FPS drops below the threshold (30) with the third players. This is not visible in Figure 1(c) because there CPU utilization is plotted for all 8 cores while there is only one main thread for all players, which runs on a single CPU core. That thread is the bottleneck. We can improve Capsule implementation to have separate main thread per player, a future work.
The same conclusion holds in Figure 1(d)–(f) and Figure 1(g)–(i), for O3DE Multiplayer Sample (O3DE, 2025a) and Exhibition, respectively: the resource footprint of additional players is sublinear. For example, as shown in Figure 1(i) for Exhibition application, one player consumes 18% of GPU while 8 players consume around 99%: a sublinear per-player increase. Similarly, CPU consumption increases sublinearly: only by up to 21% point with 8 players (28%–37%) vs. one player (16%-18%). Note that O3DE Multiplayer Sample also suffers aforementioned single-thread bottleneck. Exhibition does not suffer it and therefore can achieve over 99% GPU utilization for 8 players.
Figure 4. Scalability of Capsule as we increase number of players. Capsule and baseline could host up to 9 and 4 players, respectively, on the Exhibition application. Capsule accommodates up to 2.25x more players thanks to sublinear resources increase per added player.
Figure 4 shows average resource consumption increase as players join Exhibition application. We used Windows Task Manager to monitor resource utilization (CPU, RAM, GPU, VRAM). In the same game state, similar to the one in Figure 1, Capsule supported up to 9 players without dropping FPS below 30, while baseline sharply dropped to single-digit FPS after 4 players due to VRAM contention. At the peak of the baseline (4 players), Capsule used 3.70x less CPU, 3.87x less RAM, 1.43x less GPU, and 3.11x less VRAM compared to the baseline. Thanks to these savings, Capsule accommodates more players, beyond the baseline. As expected, baseline resource consumption increases linearly with the number of players. However, the increase in Capsule is sublinear thanks to cross-player sharing, i.e., Capsule requires a sublinear amount of extra CPU, RAM, VRAM past the first player. The same trend holds in GPU consumption as well, but since each player has a different view of the scene, sharing of GPU computation is lower than in other resources. The GPU utilization benefit of Capsule could be further improved if the application uses more shareable rendering techniques, such as shadowmaps, global illumination, and cross-view diffuse and effects sharing (Weinrauch et al., 2023).
4. Conclusion
We designed, implemented, and evaluated Capsule: an in-game-engine player isolation mechanism. Our implementation in O3DE shows that Capsule is application agnostic. We ported four existing applications to Capsule-based O3DE without application changes. Our experiments show that Capsule can increase datacenter resource–CPU, GPU, RAM, and VRAM–utilizations by accommodating up to 2.25x more players.